y(y2+1)dx+x(y2−1)dy=0


1
Expert's answer
2019-09-05T09:09:31-0400

 Separate the variables by moving all the y terms to one side of the equation and all the x terms to the other side:


"x(y^2-1)dy=-y(y^2+1)dx""(y^2-1)dy=-\frac{y(y^2+1)dx}{x}""\frac{y^2-1}{y(y^2+1)}dy=-\frac{dx}{x}"

Integrate both sides of the equation :


"\int\frac{y^2-1}{y(y^2+1)}dy=-\int\frac{dx}{x}"

Solve the first integral using partial fractions:


"\frac{y^2-1}{y(y^2+1)}=\frac{A}{y}+\frac{By+C}{y^2+1}=\frac{A(y^2+1)+(By+C)y}{y(y^2+1)}"

Then

"y^2-1=A(y^2+1)+(By+C)y""y^2-1=Ay^2+A+By^2 +Cy""y^2-1=(A+B)y^2+A+Cy""\begin{cases} A+B=1 \\ A=-1 \\ C=0 \end{cases}""\begin{cases} A=-1\\ B=2 \\C=0 \end{cases}"

Hence


"\int\frac{y^2-1}{y(y^2+1)}dy=\int(-\frac{1}{y}+\frac{2y}{y^2+1})dy="

"=-\int\frac{dy}{y}+\int\frac{d(y^2+1)}{y^2+1}=-ln y+ln (y^2+1)"

Then the solution of the differential equation is:


"-ln y+ln (y^2+1)=-\int\frac{dx}{x}""ln\frac{y^2+1}{y}=-ln{x}+ln{C}""ln\frac{y^2+1}{y}=ln\frac{C}{x}""\frac{y^2+1}{y}=\frac{C}{x}"

Solve for y:


"y^2+1=\frac{C}{x}y""y^2-\frac{C}{x}y+1=0"


"y=\frac{\frac{C}{x}\pm\sqrt{\frac{C^2}{x^2}-4}}{2}""y=\frac{\frac{C}{x}\pm\frac{\sqrt{C^2-4x^2}}{x}}{2}""y=\frac{C\pm\sqrt{C^2-4x^2}}{2x}"

Answer: "y=\frac{C\pm\sqrt{C^2-4x^2}}{2x}"